OpenStudy (anonymous):

]Choose the equation below that represents the line passing through the point (1, -4) with a slope of 1/2... i beg you plllzzzzzzzzz help :((

OpenStudy (anonymous):

\[m=\frac{1}{2}\] \[(x1,y1)=(1, -4) \]

OpenStudy (anonymous):

You have given the slope and point.. Use the following: \[\large y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

i dont get it omggggg :(

OpenStudy (anonymous):

what don't you get you have to be specific.

OpenStudy (anonymous):

\[y + 4 = \frac{1}{2}(x - 1)\] Solve this..

OpenStudy (anonymous):

the whole question

OpenStudy (anonymous):

there you got the answer

OpenStudy (anonymous):

ooo now i get ittt

OpenStudy (anonymous):

that is the answer

OpenStudy (anonymous):

y+4=1/2(x-1) right??

OpenStudy (anonymous):

Just like that you got it?

OpenStudy (anonymous):

thanks romero